### v1

# A quantum Monte Carlo algorithm for arbitrary high-spin Hamiltonians

Abstract
We present a universal parameter-free quantum Monte Carlo algorithm for simulating arbitrary high-spin (spin greater than $1/2$) Hamiltonians. This approach extends a previously developed method by the authors for spin-$1/2$ Hamiltonians [Phys. Rev. Research 6, 013281 (2024)]. To demonstrate its applicability and versatility, we apply our method to the spin-$1$ and spin-$3/2$ quantum Heisenberg models on the square lattice. Additionally, we detail how the approach naturally extends to general Hamiltonians involving mixtures of particle species, including bosons and fermions. We have made our program code freely accessible on GitHub.
I Introduction
Quantum high-spin models, namely quantum many-body Hamiltonians depicting the interactions between particles possessing spins greater than $1/2$, are pivotal for understanding complex magnetic interactions, quantum phase transitions, topological phases, and quantum entanglement in condensed matter systems [1, 2, 3]. Such models provide a rich framework for both theoretical studies and experimental realizations, enabling the exploration of new physical phenomena that are not present in simpler spin-$1/2$ models including quantum phase transitions, topological order, entanglement, and exotic magnetic states. Some notable examples are the high-spin Heisenberg models, Haldane chains which include an exchange interaction and a single-ion anisotropy term, and the Affleck-Kennedy-Lieb-Tasaki (AKLT) model, which serves as an example of a spin-1 chain with a ground state that exhibits a Haldane gap and non-trivial topological order [4].
Despite the importance of high-spin models in condensed matter physics, numerical simulation methods for studying these systems are generally scarce (for exceptions addressing specific models, see Refs. [5, 6, 7]). Specifically, quantum Monte Carlo (QMC) techniques, which are essentially the only viable method for studying the finite-temperature equilibrium properties of large-scale quantum many-body systems without approximations, have not yet been fully adapted to the simulation of high-spin quantum systems.
This work aims to remedy this situation by providing a universal framework for studying arbitrarily complex high-spin quantum models of any spin value and interactions of essentially any geometry, dimension, locality, and connectivity. Furthermore, a natural generalization of the proposed framework to address QMC simulations of mixed-species models including spin-$1/2$ particles, bosons, and fermions is also discussed in detail.
The present technique builds on a recent study by the authors [8] in which a universal QMC algorithm designed to simulate arbitrary spin-$1/2$ Hamiltonians was devised. There, it was shown that casting the to-be-simulated Hamiltonian in Permutation Matrix Representation (PMR) form [9] allows one to write the partition function of any spin-$1/2$ system as a sum of efficiently calculable terms each of which is associated with a closed walk on the Hamiltonian graph and where Monte Carlo updates to faithfully sample these walks may be automatically generated in a systematic way [9, 8].
In this paper, we generalize the aforementioned technique to the case of high-spin (spin greater than $1/2$) Hamiltonians and introduce a similar-in-spirit universal Monte Carlo algorithm designed to reliably simulate arbitrary high-spin systems. To that aim, we devise a protocol for generating the necessary set of QMC updates, based on PMR decomposition, that ensure an ergodic Markov chain Monte Carlo sampling of the partition function of essentially any conceivable input system. We illustrate that while for spin-$1/2$ systems achieving the same goal required finding the nullspace of sets of binary (modulo 2) vectors representing the permutation matrices of the Hamiltonian, for spin-$s$ particles the task is generalized to finding a similar set of modulo $(2s+1)$ vectors.
The paper is organized as follows. In Sec. II, we provide a brief overview of the permutation matrix representation quantum Monte Carlo (PMR-QMC) and analyze high-spin Hamiltonians in this context. In Sec. III, we discuss the QMC algorithm, describing in detail the method we have devised to generate all the necessary QMC updates and proving that these moves ensure both ergodicity and detailed balance. In Sec. IV we showcase the power of our technique by presenting simulation results for two models, namely, the spin-$1$ and spin-$3/2$ quantum Heisenberg models on the square lattice, and in Sec. V we discuss in detail how the approach taken can be extended to include other particle species as well as mixtures thereof. We conclude in Sec. VI with an additional discussion and future directions of research.
II Permutation matrix representation for high-spin Hamiltonians
II.1 Overview of PMR-QMC
We begin by providing a brief overview of the permutation matrix representation (PMR) protocol [9, 8] on which the simulation algorithm will be based. PMR begins by first casting the to-be-simulated Hamiltonian in PMR form, i.e., as a sum
$H=D_{0}+\sum_{j=1}^{M}D_{j}P_{j}\,,$ (1)
where $D_{j}$ are diagonal matrices and $\{P_{j}\}_{j=1}^{M}$ are permutation matrices that are a subset of a special Abelian group, $G$  [^1] Every $P\in G$ is a permutation matrix with no nonzero diagonal elements, except for the identity matrix $P_{0}={\mathbb{1}}$. For every $|i\rangle,|j\rangle$ pair, there exists a unique $P\in G$ such that $P|i\rangle=|j\rangle$. See [15] for more details.. For spin-1/2 systems, $G$ consists of all Pauli-X strings [8].
A non-trivial, though useful, consequence is that we can write the partition function as a sum of ‘generalized Boltzmann weights’ [9, 11, 8]
$\mathcal{Z}=\sum_{z}\sum_{q=0}^{\infty}\sum_{S_{{\bf{i}}_{q}}=\mathbb{1}}D_{(z,S_{{\bf{i}}_{q}})}{{e}}^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}=\sum_{\cal C}w_{\mathcal{C}},$ (2)
where each $\mathcal{C}=(z,S_{{\bf{i}}_{q}})$ is a QMC configuration, and each weight
$w_{\cal C}=D_{(z,S_{{\bf{i}}_{q}})}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}$ (3)
[…]
To construct a Markov chain, it is necessary to employ real-valued non-negative weights, whether or not a sign problem [14] is present. To this end, one can consider either absolute values of real components of the weights or absolute values of the weights:
$W_{{\cal C}}^{(1)}=\left|\operatorname{Re}[w_{\cal C}]\right|,\quad W_{{\cal C}}^{(2)}=\left|w_{\cal C}\right|\,.$ (4)
The PMR formulation allows one to measure a wide range of static operators and dynamical quantities [15]. The key to being able to do so is to write for any given operator $A$ its thermal average as
$\langle A\rangle=\frac{{\mbox{Tr\,}}[A{{e}}^{-\beta H}]}{{\mbox{Tr\,}}[{{e}}^{-\beta H}]}=\frac{\sum_{\cal C}A_{\cal C}w_{\cal C}}{\sum_{\cal C}w_{\cal C}}.$ (5)
Although, generally, both $w_{\cal C}$ and $A_{\cal C}$ are complex-valued, both sums $\sum_{\cal C}A_{\cal C}w_{\cal C}$ and $\sum_{\cal C}w_{\cal C}$ are real-valued since both $H$ and $A$ are Hermitian operators. Therefore, we have
$\langle A\rangle$ $=$ $\frac{\sum_{\cal C}\left(\operatorname{Re}[A_{\cal C}w_{\cal C}]/W_{\cal C}^{(1)}\right)\cdot W_{\cal C}^{(1)}}{\langle\textrm{sgn}\rangle_{1}\sum_{\cal C}W_{\cal C}^{(1)}},$ (6)
$\langle A\rangle$ $=$ $\frac{\sum_{\cal C}\left|A_{\cal C}\right|\cos(\arg(A_{\cal C}w_{\cal C}))\cdot W_{\cal C}^{(2)}}{\langle\textrm{sgn}\rangle_{2}\sum_{\cal C}W_{\cal C}^{(2)}},$ (7)
where $\textrm{sgn}_{1}(w_{\cal C})=\textrm{sgn}(\operatorname{Re}[w_{\cal C}])$, $\textrm{sgn}_{2}(w_{\cal C})=\cos(\arg(w_{\cal C}))$, and
$\langle\textrm{sgn}\rangle_{1}$ $=$ $\frac{\sum_{\cal C}\textrm{sgn}_{1}(w_{C})\cdot W_{\cal C}^{(1)}}{\sum_{C}W_{\cal C}^{(1)}},$ (8)
$\langle\textrm{sgn}\rangle_{2}$ $=$ $\frac{\sum_{\cal C}\textrm{sgn}_{2}(w_{C})\cdot W_{\cal C}^{(2)}}{\sum_{C}W_{\cal C}^{(2)}}.$ (9)
The values of $\operatorname{Re}[A_{\cal C}w_{\cal C}]/W_{\cal C}^{(1)}$ and $\left|A_{\cal C}\right|\cos(\arg(A_{\cal C}w_{\cal C}))$, therefore, represent the instantaneous quantity associated with the configuration ${\cal C}={(z,{S_{{\bf{i}}_{q}}})}$ that will be collected during the simulation when using the weights $W_{\cal C}^{(1)}$ and $W_{\cal C}^{(2)}$, respectively. For further details about calculating a wide range of observables throughout the simulation, see Refs. [9, 8, 15].
II.2 PMR decomposition of single spin operators
Before discussing the PMR formulation of general high-spin Hamiltonians, let us first briefly review high-spin operators. The matrix elements of the $(2s+1)\times(2s+1)$ spin matrices $X$, $Y$, and $Z$ for $s\in\{1/2,1,3/2,2,5/2,\ldots\}$ are given by
$X_{jk}$ $=\frac{1}{2}\left(\delta_{j,k+1}+\delta_{j+1,k}\right)\sqrt{(s+1)(j+k-1)-jk},$
$Y_{jk}$ $=\frac{j}{2}\left(\delta_{j,k+1}-\delta_{j+1,k}\right)\sqrt{(s+1)(j+k-1)-jk},$
$Z_{jk}$ $=(s+1-j)\delta_{j,k},$
where $j,k\in\{1,2,\dots,2s+1\}$. In the spirit of PMR, these matrices can be written in the form
$X$ $=D^{+}P+D^{-}P^{-1},$ (10)
$Y$ $=-iD^{+}P+iD^{-}P^{-1},$ (11)
$Z$ $=D^{(z)}\,,$ (12)
respectively, where $P$ is the $(2s+1)\times(2s+1)$ permutation matrix given by
$P=\begin{pmatrix}0&1&0&\ldots&0\\ 0&0&1&\ldots&0\\ \vdots&\vdots&\ddots&\ddots&0\\ 0&0&\ldots&0&1\\ 1&0&\ldots&0&0\\ \end{pmatrix},$ (13)
which obeys $P^{2s+1}=\mathbb{1}$, and $D^{+}$, $D^{-}$, $D^{(z)}$ are diagonal matrices with the following diagonal entries:
$D^{(z)}_{j}$ $=s+1-j,\,$ (14)
$D^{+}_{j}$ $=\frac{1}{2}\,\sqrt{(2s-j+1)j},\,$ (15)
$D^{-}_{j}$ $=\frac{1}{2}\,\sqrt{(2(s+1)-j)(j-1)},$ (16)
with $j=1,\ldots,2s+1$.
For reasons that will become clear shortly, we also introduce the matrices $D^{(k)}$ and $D^{(z,k)}$ such that
$D^{(k)}_{j}$ $=$ $D^{+}_{(j+k)\bmod{(2s+1)}},$ (17)
$D^{(z,k)}_{j}$ $=$ $D^{(z)}_{(j+k)\bmod{(2s+1)}},$ (18)
where $D^{+}_{0}$ and $D^{(z)}_{0}$ are defined as $D^{+}_{0}=D^{+}_{2s+1}=0$ and $D^{(z)}_{0}=D^{(z)}_{2s+1}=-s$. We note that $D^{+}=D^{(0)}$, $D^{-}=D^{(-1)}$, and $D^{(z)}=D^{(z,0)}$. The matrices $D^{(k)}$ and $D^{(z,k)}$ obey
$PD^{(k)}$ $=$ $D^{(k+1)}P,$ (19)
$P^{-1}D^{(k)}$ $=$ $D^{(k-1)}P^{-1},$ (20)
$PD^{(z,k)}$ $=$ $D^{(z,k+1)}P,$ (21)
$P^{-1}D^{(z,k)}$ $=$ $D^{(z,k-1)}P^{-1}.$ (22)
II.3 High-spin Hamiltonians
Consider now an $n$-particle Hamiltonian given as the linear combination
$H=\sum_{i}c_{i}\prod_{k=1}^{m_{i}}s^{(i)}_{j_{i,k}},$ (23)
where $c_{i}$ are real-valued coefficients and $\prod_{k=1}^{m_{i}}s^{(i)}_{j_{i,k}}$ are spin operator strings. Here, $s^{(i)}_{j_{i,k}}$ represents a spin matrix $s\in\{X,Y,Z\}$ in the $i$-th string acting on the $j_{i,k}$-th particle, where $j_{i,k}\in\{1,2,\dots,n\}$. The operator $s^{(i)}_{j_{i,k}}$ is a tensor product of a spin matrix and $n-1$ identity matrices such that it has the same matrix dimension as the Hamiltonian. Each spin operator string may contain any number of any of the three spin matrices for each of the particles in any order.
Equation (23) can be rewritten as
$H=\sum_{i}c_{i}\bigotimes_{j=1}^{n}\prod_{k=1}^{m_{i,j}}s^{(i)}_{j,k},$ (24)
where $s^{(i)}_{j,k}\in\{X,Y,Z\}$ denotes a spin matrix in the $i$-th term acting on the $j$-th particle, and $m_{i,j}$ are non-negative integers.
We will now cast the general Hamiltonian $H$, Eq. (24), in PMR form. To do that, we apply Eqs. (10)–(12) to each $s^{(i)}_{j,k}$ and rewrite the Hamiltonian in the form
$H=\sum_{i}d_{i}\bigotimes_{j=1}^{n}\prod_{k=1}^{m_{i,j}}t^{(i)}_{j,k},$ (25)
where each $t^{(i)}_{j,k}$ acts on the $j$-th particle and is either $D^{(0)}P$ or $D^{(-1)}P^{-1}$ or $D^{(z,0)}$.
Next, we ‘push’ all diagonal matrices in each product $\prod_{k=1}^{m_{i,j}}t^{(i)}_{j,k}$ to the left using the relations (19)–(22) to obtain
$H=\sum_{i}d_{i}\bigotimes_{j=1}^{n}D^{(i,j)}\bigotimes_{j=1}^{n}P^{(i,j)},$ (26)
where $D^{(i,j)}$ and $P^{(i,j)}$ are diagonal and permutation matrices, respectively, and the index $j$ indicates the action on the $j$-th particle. The matrix $P^{(i,j)}$ is equal to the matrix $P$ raised to some power $n_{i,j}\in\{0,1,\dots,2s\}$.
Last, we group together all terms that have the same $\bigotimes_{j=1}^{n}P^{(i,j)}$ component, ending up with a Hamiltonian of the form
$H=\sum_{i}D_{i}P_{i},$ (27)
where $P_{i}=\bigotimes_{j=1}^{n}P^{n_{i,j}}$.
We have thus achieved a PMR decomposition for arbitrary spin-$s$ Hamiltonians.
III The QMC algorithm
III.1 QMC configurations
For any Hamiltonian cast in PMR form, the partition function $Z={\mbox{Tr\,}}[e^{-\beta H}]$ can be written as a sum of configuration weights [cf. Eq. (2)], where a configuration ${\cal C}=\{|z\rangle,S_{{\bf{i}}_{q}}\}$ is a pair of a classical (diagonal) basis state $|z\rangle$ and a product $S_{{\bf{i}}_{q}}$ of permutation operators that must evaluate to the identity element $P_{0}=\mathbb{1}$. The configuration ${\cal C}$ induces a list of states $\{|z_{0}\rangle=|z\rangle,|z_{1}\rangle,\ldots,|z_{q}\rangle=|z\rangle\}$, which in turn generates a corresponding multi-set of energies $E_{{\cal C}}=\{E_{z_{0}},E_{z_{1}},\ldots,E_{z_{q}}\}$ for the configuration.
We can now consider a QMC algorithm that samples these configurations with probabilities proportional to their weights $W_{{\cal C}}$, Eq. (4). The Markov process would start with some initial configuration and a set of (probabilistic) rules, or QMC updates, will dictate transitions from one configuration to the next.
We will take the initial state to be $\mathcal{C}_{0}=\{|z\rangle,S_{0}=\mathbb{1}\}$ where $|z\rangle$ is a randomly generated initial classical state. The weight of this initial configuration is
$W_{{\cal C}_{0}}={{e}}^{-\beta[E_{z}]}={{e}}^{-\beta E_{z}}\,,$ (28)
i.e., the classical Boltzmann weight of the initial randomly generated basis state $|z\rangle$.
The set of required QMC updates will be discussed in the next sections. To ensure that configurations are sampled properly, i.e. in proportion to their weight, one must ensure that the Markov process is ergodic, i.e., that the QMC updates are capable of generating all basis states $|z\rangle$ as well as all sequences $S_{{\bf{i}}_{q}}$ evaluating to the identity. An additional sufficient requirement to ensure proper sampling is that of detailed balance, which dictates that the ratio of transition probabilities from one configuration to another and the transition in the opposite direction equals to the ratio of their respective weights [16, 17]. In the following, we show how both conditions are made to be satisfied for general high-spin Hamiltonians with arbitrary interactions.
III.2 Fundamental cycles
It follows from Eqs. (26) and (27) that the permutation operators $P_{i}$ are tensor products of powers of $P$. As such, (i) all permutation operators commute, and (ii) $P^{2s+1}=\mathbb{1}$, so each permutation operator $P_{i}=\bigotimes_{j=1}^{n}P^{n_{i,j}}$ satisfies $P_{i}^{2s+1}=\mathbb{1}$.
Denoting by $p_{i}$ the integer-string $[n_{i,1}n_{i,2}\cdots n_{i,n}]$, one can easily verify that the product of two permutation operators $P_{i}$ and $P_{k}$ would likewise correspond to modular addition of the components of $p_{i}$ and $p_{k}$ modulo $12s+$ 1, i.e.
$P_{i}P_{k}\rightarrow p_{i}+p_{j}\mod 2s+1\,.$ (29)
For a sequence of operators evaluating to the identity, $S_{{\bf{i}}_{q}}=P_{i_{q}}\ldots P_{i_{1}}$, we have $\sum_{j=1}^{q}p_{i_{j}}\equiv{\bf 0}\,(\text{mod }2s+1)$, where ${\bf 0}$ is an integer-string consisting of only zeros. We note that $S_{{\bf{i}}_{q}}$ is a permutation of the multiset of operators $\{P_{i_{1}},\dots,P_{i_{q}}\}$, which can be represented as an integer string $[a_{1}\,a_{2}\,\dots\,a_{M}]$, where $a_{k}$ denotes the number of occurrences of the operator $P_{k}$ among $P_{i_{1}},\dots,P_{i_{q}}$.
The question of how one can generate all possible sequences of operators (which evaluate to the identity) is therefore reduced to the question of how one can generate all integer-strings $[a_{1}\,a_{2}\,\dots\,a_{M}]$ that obey the following system of linear equations over mod-$(2s+1)$ addition
$\sum_{i=1}^{M}a_{i}\cdot n_{i,j}\equiv 0\pmod{2s+1},\,j=1,2,\dots,n.$ (30)
Equation (30) can be solved by finding the null space basis over addition modulo $12s+$ 1 for the matrix ${\bf n}^{T}$ whose columns are $p_{1}^{T},\ldots,p_{M}^{T}$. For details on accomplishing this task regardless of the decomposition of $12s+$ 1 into prime factors, the reader is referred to Appendix B.
We shall call a multiset of permutation operators that multiply to the identity a ‘cycle’. The length of a cycle would be the number of distinct permutation operators in it. We shall refer to cycles represented by the integer strings from the null space basis as fundamental cycles.
Generally, the nullspace basis states can be chosen in many different ways, and so the choice of the set of fundamental cycles is not unique. From a practical point of view, however, we find that obtaining a ‘minimal cycle basis’, i.e., a basis that minimizes the lengths of all basis cycles, is advantageous from the QMC standpoint. This follows from the fact that the probability of a QMC update to be accepted is a decreasing function of the cycle length. To reduce the cycles lengths, we therefore find a null space basis and then proceed to replace long-cycle basis states with shorter basis states by performing mod-$(2s+1)$ additions between the integer-strings of pairs of cycles, accepting the changes each time a new cycle with a shorter length is found. The process ends when a pass through all pairs of cycles does not result in an improvement.
It follows from the properties of null space that any cycle can be obtained by (i) insertion and removal of fundamental cycles, (ii) insertion and removal of trivial cycles consisting of $12s+$ 1 identical permutation operators, and (iii) swapping the order of two adjacent permutation operators.
Update Spin-$1/2$ QMC simulations High-spin QMC simulations
Pair insertion and deletion The pair consists of two identical permutation operators. The pair consists of two mutually inverse permutation operators, which can be either the same or different.
Classical update The update performs a spin-flip of a random spin. The update selects a random particle and changes its spin value to a randomly selected different value.
Fundamental cycle completion The update that consists of choosing a subsequence $S$ from $S_{{\bf{i}}_{q}}$, choosing a fundamental cycle containing all operators of the subsequence $S$, and attempting to replace the subsequence $S$ with the remaining operators from the selected cycle. A subsequence $S$ must not contain repeated operators. The elements of $S$ are not required to be consecutive within $S_{{\mathbf{i}}_{q}}$. If $A$ and $B$ are two complementary parts of a cycle, we do not directly replace $A$ with $B$. Instead, we replace either $A^{-1}$ by $B$ or $A$ by $B^{-1}$. The selected subsequence $S$ in $S_{{\mathbf{i}}_{q}}$ may now contain repetitions, and the expressions for the acceptance probability are adjusted accordingly. See details in Appendix A.
Worm update The update is rejected with a small probability $p_{f}$ at each intermediate step, where $p_{f}$ is an adjustable parameter. Intermediate configurations are assigned their ‘natural’ weight $W_{\cal C}$ as per Eq. (4). We find it useful not to artificially exit the worm, i.e., we use $p_{f}=0$. To prevent the worm from straying too far from being healed, each intermediate configuration $\cal C$ is assigned the weight $W_{\cal C}\exp(-\alpha d)$. where $W_{\cal C}$ is defined by Eq. (4), $\alpha$ is an adjustable parameter that may depend on system size and temperature, and $d$ is the ‘distance from identity,’ defined as the number of particles with different spin values in $|z_{0}\rangle$ and $|z_{q}\rangle$.
Table 1: Main differences in QMC updates between the spin-$1/2$ case (covered in Ref. [8]) and the high-spin case considered in the present work.
III.3 QMC updates
As a preliminary step, prior to the simulation taking place, we carry out the PMR decomposition of the Hamiltonian, and produce a list of fundamental cycles for the to-be-simulated Hamiltonian cast in PMR form (see details in Sections above). Additionally, we add $M$ trivial cycles to the list of fundamental cycles, each containing $12s+$ 1 identical permutation operators.
Because the Hamiltonian is Hermitian, for each operator $P_{i}$ from the PMR decomposition (1), the operator $P_{i}^{-1}$ is also included in the PMR. Hence, the list of permutation operators consists of pairs of mutually inverse operators, as well as the operators that are inverse to themselves.
We employ the following QMC updates: (i) the worm update, (ii) block swap, and (iii) classical updates.
The worm update involves single operator moves performing a ‘disturbance’ causing $S_{{{\bf i}}_{q}}$ to evaluate to a non-identity permutation, as well as ‘healing’ back to an identity-forming sequence (see details in Ref. [8]). Within the worm update, we also employ local swaps, pair insertions and deletions, and fundamental cycle completions.
The basic QMC updates mentioned above are similar in nature to analogous moves used in the spin-$1/2$ case [8], with a few necessary important changes which are summarized in Table 1. In particular, fulfillment of the detailed balance condition can be shown similarly to the spin-$1/2$ case [8].
It was shown in the previous section that local swaps, fundamental cycle completions, and trivial cycle completions are sufficient to ensure ergodicity along the quantum (or imaginary-time) dimension, i.e., the ability to generate all permutation operator sequences that evaluate to the identity. Also, we find applying the worm update very useful in practice because it further accelerates Markov chain mixing and the achievement of ergodicity in the quantum dimension.
The ergodicity along the classical dimension, i.e., the generation of all possible classical basis states $|z\rangle$, is achieved by employing the classical update.
For two arbitrary configurations ${\cal C}=\{|z\rangle,S_{{\bf{i}}_{q}}\}$ and ${\cal C^{\prime}}=\{|z^{\prime}\rangle,S_{{\bf{i^{\prime}}}_{q^{\prime}}}\}$ such that $W_{\cal C}\neq 0$ and $W_{\cal C^{\prime}}\neq 0$, the above QMC updates allow in particular the following sequence of transformations: ${\cal C}\to{\cal C}_{0}\to{\cal C}^{\prime}_{0}\to{\cal C}^{\prime}$, where ${\cal C}_{0}=\{|z\rangle,\mathbb{1}\}$ and ${\cal C}^{\prime}_{0}=\{|z^{\prime}\rangle,\mathbb{1}\}$. Therefore, the transformation from ${\cal C}$ to ${\cal C}^{\prime}$ is possible, and the ergodicity holds in the entire configuration space.
Figure 1: Calculations of the spin-$1$ quantum Heisenberg model on a square $L\times L$ lattice. Top: Specific heat as a function of inverse-temperature $\beta$. Bottom: magnetic susceptibility as a function of $\beta$. The standard errors are no larger than the size of a marker.
Figure 2: Calculations of the spin-$3/2$ quantum Heisenberg model on a square $L\times L$ lattice. Top: Specific heat as a function of inverse-temperature $\beta$. Bottom: magnetic susceptibility as a function of $\beta$. The standard errors are no larger than the size of a marker.
IV Simulating high-spin quantum Heisenberg models
In this section, we illustrate the effectiveness of our method by probing the thermodynamic behavior of the high-spin quantum Heisenberg model on a square $L\times L$ lattice with open boundary conditions under an external magnetic field. The Hamiltonian of the model is given by
$H=-J\sum_{\langle i,j\rangle}\left(X_{i}X_{j}+Y_{i}Y_{j}+Z_{i}Z_{j}\right)+h\sum_{i=1}^{n}Z_{i}\,,$ (31)
where $n=L^{2}$ is the number of particles, $\langle i,j\rangle$ denotes neighbors on the lattice, and we choose for our simulations the interaction strength to be $J=1$ and the external magnetic field strength to be $h=0.1$.
For these chosen parameters, we compute via our QMC algorithm the specific heat
$C=\beta^{2}\left(\langle E^{2}\rangle-\langle E\rangle^{2}\right),$ (32)
and the magnetic susceptibility
$\chi_{M}=\beta\left(\langle M^{2}\rangle-\langle M\rangle^{2}\right),$ (33)
where above $M$ is the $Z$-magnetization $M=\frac{1}{sN}\sum_{i=1}^{N}Z_{i}$. We have computed the dependence of both quantities on the inverse temperature $\beta$ across various system sizes.
In Fig. 1 the results for a system of spin-$1$ particles is presented, whereas Fig. 2 displays data for systems of particles with spin values of $s=3/2$. An anticipated feature is the shift in the peak positions of both specific heat and magnetic susceptibility as the system size increases. The peaks become sharper and shift closer to the critical temperature in the thermodynamic limit, indicating finite-size corrections.
The peaks for $s=3/2$ appear at smaller $\beta$ values compared to $s=1$. This is expected since the larger energy level separations in higher-spin systems increase their resistance to thermal excitations, allowing ordered states to persist at higher critical temperatures.
In addition, higher-spin particles are expected to exhibit reduced quantum fluctuations, leading to more classical behavior at low temperatures. The peaks for spin-$3/2$ appear more well-defined compared to the broader peaks in spin-$1$ systems, suggesting that quantum fluctuations smear out the critical behavior over a wider temperature range.
The above numerical results highlight the rich thermodynamic behavior of high-spin Heisenberg models, emphasizing the interplay between system size, spin magnitude, and thermal fluctuations.
The proposed QMC algorithm can also be readily applied to other high-spin Heisenberg models—whether on the square lattice or on any other graph.
V Extending the framework to arbitrary mixed Hamiltonians
The method presented above can be extended to ’mixed-spin’ Hamiltonians, where particles of different species exist and interact. To generalize the method for this case, each particle is assigned a spin value $s_{i}$, rather than a single global spin value $s$ shared by all particles as in the previous derivations.
Moreover, incorporating fermionic and bosonic degrees of freedom can be readily accomplished, as we outline below.
V.1 Mixed-spin Hamiltonians
For a mixed-spin Hamiltonian, the permutation operators can be written as
$P_{i}=\bigotimes_{j=1}^{n}P^{(i,j)}=\bigotimes_{j=1}^{n}P(2s_{j}+1)^{n_{i,j}},$ (34)
where $P(2s_{j}+1)$ is the $(2s_{j}+1)\times(2s_{j}+1)$ permutation matrix given by Eq. (13). As a consequence, each permutation operator $P_{i}$ can be represented as an integer-string $p_{i}=[n_{i,1}n_{i,2}\cdots n_{i,n}]$, where $n_{i,j}\in\{0,1,\ldots,2s_{j}\}$ (cf. Sec. III.2).
Similar to the single species case, a multiset of operators $\{P_{i_{1}},\dots,P_{i_{q}}\}$ is represented by an integer string $[a_{1}\,a_{2}\,\dots\,a_{M}]$, where $a_{k}$ denotes the number of occurrences of the operator $P_{k}$ among $P_{i_{1}},\dots,P_{i_{q}}$. In order to find fundamental cycles for the to-be-simulated Hamiltonian, one needs to find integer strings $[a_{1}\,a_{2}\,\dots\,a_{M}]$ such that
$\sum_{i=1}^{M}a_{i}\cdot n_{i,j}\equiv 0\,\pmod{2s_{j}+1},\,j=1,2,\dots,n\,.$ (35)
For $s_{1}=s_{2}=\ldots=s_{N}=s$, the above problem reduces to Eq. (30). The more general case may also be solved efficiently. We refer the reader to Appendix B for details on how to solve Eq. (35).
V.2 Incorporation of fermions
The addition of fermions to any mixed-spin model may be accomplished by converting the fermionic degrees of freedom to a spin-$1/2$ particle representation. This is carried out via the application of a Jordan-Wigner transformation (JWT) [18] which maps the second-quantized annihilation operator $c_{j}$ to an operator on $j$ spins according to
$c_{j}\to\left(\prod_{k=1}^{j-1}Z_{k}\right)\frac{X_{j}-iY_{j}}{2}$ (36)
so that $c^{\dagger}_{j}c_{j}=({\mathbb{1}}+Z_{j})/2$. To write the Fermi-Hubbard Hamiltonian in PMR form, we rewrite the JWT as products of a diagonal operator (a function of Pauli-$Z$ strings) and a permutation operator (a Pauli-$X$):
$c_{j}\to\left[\left(\prod_{k=1}^{j-1}Z_{k}\right)\frac{{\mathbb{1}}-Z_{j}}{2}\right]X_{j}\,,$ (37)
$c_{j}^{\dagger}\to\left[\left(\prod_{k=1}^{j-1}Z_{k}\right)\frac{{\mathbb{1}}+Z_{j}}{2}\right]X_{j}\,.$ (38)
Once the fermionic sites are labeled, the fermionic degrees of freedom are mapped to spin-half operators, in which case the Hamiltonian reverts to being a mixed-spin model again.
V.3 Incorporation of bosons
As for the inclusion of bosonic degrees of freedom, we will see in this section that the creation and annihilation operators in the second-quantized basis correspond to permutation operators of infinite order (see also [19]).
As the computational basis for the PMR expansion, we use the second quantized occupation number basis for bosons, where a basis state is given as $|{\bf n}\rangle=|n_{1},n_{2},\ldots,n_{L}\rangle$ with $L$ being the number of sites and $n_{1},\ldots,n_{L}$ are nonnegative integers representing the number of bosons in each site. We denote the total number of bosons, $\sum_{i=1}^{L}n_{i}$, by $n$. The operators $\hat{b}_{i}^{{\dagger}},\hat{b}_{i}$ are creation and annihilation operators, respectively, obeying
$\hat{b}_{i}^{{\dagger}}|n_{1},\ldots,n_{i},\ldots,n_{L}\rangle=\sqrt{(n_{i}+1)}|n_{1},\ldots n_{i+1},\ldots,n_{L}\rangle\,,$ (39)
where $|{\bf n}^{(i,j)}\rangle$ stands for the state $|{\bf n}\rangle$ with one additional boson at site $i$ and one fewer at site $j$. The operator $\hat{n}_{i}=\hat{b}_{i}^{{\dagger}}\hat{b}_{i}$ is the number operator. Define the following operator on the state in Fock space $|{\bf{n}}\rangle=|n_{1},\ldots,n_{i},\ldots,n_{j},\ldots,n_{L}\rangle$
$P_{m}|{\bf{n}}\rangle$ $=|{\bf{n}}^{(i_{m},j_{m})}\rangle$
$\equiv|n_{1},\ldots,n_{i_{m}}+1,\ldots,n_{j_{m}}-1,\ldots,n_{L}\rangle\,.$ (40)
$P_{m}$ can be thought of as a permutation operator, permuting the states with different number of bosons on $i_{m}$ and $j_{m}$ site. It is important to note that application of $b^{\dagger}_{i_{m}}b_{j_{m}}$ would annihilate a state with $n_{j_{m}}=0$. However, one can allow $P_{m}$ to simply map states with $n_{j_{m}}=0$ to states with negative $n_{j_{m}}$. The annihilation will be imposed by the matrices $D_{m}$.
Given this construction, for every $P_{m}$ one can write a corresponding inverse permutation $(P_{m})^{-1}$ that reverses the mapping. Moreover, $P_{m}$ loses its finite periodicity, as there is no longer any finite $s$ such that $P_{m}^{2s+1}={\mathbb{1}}$. Thus, the permutation operators $P_{m}$ have infinite order. In this case, finding identity equivalent string of permutations will be tantamount to finding the mod-$\infty$ nullspace of vector of integer. This is nothing but solving a system of linear equations with integer solutions, i.e. a system of Diophantine equations.
Given a set of bosonic permutations $\{P_{i}\}$, we can equivalently use integer strings $\{\boldsymbol{p}_{i}\}$ to denote the action of a particular $P_{i}$ on the a Fock space basis state via an integer string $\boldsymbol{p}_{i}$. In this way, in order to find all identity equivalent string $\Pi_{i}P_{i}={\mathbb{1}}$, we will need to find the set of solutions to the following equation
$\sum_{i}a_{i}\boldsymbol{p}_{i}=\boldsymbol{0}\,,$ (41)
where $a_{i}$ are unknown integers. Equation (41) can be addressed using standard methods for Diophantine equations (see Appendix B).
VI Summary and discussion
We presented a universal, parameter-free, Trotter-error-free quantum Monte Carlo scheme capable of simulating, for the first time, arbitrary high-spin Hamiltonians. We have demonstrated that the permutation matrix representation of Hamiltonians allows one to automatically produce QMC updates that are provably ergodic and satisfy detailed balance, thereby ensuring the convergence of the Markov chain to the equilibrium.
Our algorithm therefore enables one to study the equilibrium properties of essentially any conceivable high-spin system using a single piece of code. We illustrated this ability by producing results for the quantum Heisenberg model for two types of spin particles, namely, spin-$1$ and spin-$3/2$.
We have further shown that the methodologies devised here can be extended to other particle species and mixtures thereof.
We hope that the feasibility of such calculations and applicability of our code to a wide range of high-spin systems will contribute to the broader understanding of quantum magnetism and phase transitions in low-dimensional quantum systems.
We note that while our approach guarantees a correct equilibrium distribution of the Markov chain, the algorithm does not guarantee a universal rapid mixing of the Markov chain, nor does it resolve or aim to resolve the sign problem.
We believe that the generality and versatility of the method developed here will make our proposed technique a very useful tool for condensed matter physicists studying spin systems, allowing the community to explore with ease physical models that have so far been inaccessible, cumbersome to code, or too large to implement with existing techniques. To that aim, we have made our program code freely accessible on GitHub [20].
Acknowledgements.
This project was supported in part by NSF award #2210374. The authors acknowledge the Center for Advanced Research Computing (CARC) at the University of Southern California for providing the computing resources used in this work. In addition, this material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0063. All material, except scientific articles or papers published in scientific journals, must, in addition to any notices or disclaimers by the Contractor, also contain the following disclaimer: Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA).
Appendix A Cycle completion with gaps
As was mentioned in the main text, a simple cycle completion protocol, in which consecutive elements of $S_{{\bf{i}}_{q}}$ form a subsequence $S$, may lead to a violation of the ergodicity condition and therefore also an incorrect calculation. The reason for this is the possibility of the resultant configuration having zero weight [as per Eq. (4)]. As a consequence, the fundamental cycle completion routine may never be applied for some of the fundamental cycles during the Markov process, which can in turn lead to the ergodicity violation.
To resolve this issue, we have developed a protocol that we call ‘cycle completion with gaps’. This protocol does not require the elements of the sequence $S$ to form a consecutive unit within $S_{{\bf{i}}_{q}}$. A detailed description of this subroutine follows.
The parameters are: $r_{\min}$, $r_{\max}$, $l_{\min}(r)$, $l_{\max}(r)$. One can choose: $r_{\min}=(f_{\min}-1)/2$, $r_{\max}=(f_{\max}+1)/2$, $l_{\min}(r)=2r-1$, $l_{\max}(r)=2r+1$, where $f_{\min}$ and $f_{\max}$ are minimal and maximal fundamental cycles lengths, respectively. The other reasonable option (‘exhaustive search’) is to choose $r_{\min}=0$, $r_{\max}=f_{\max}$, $l_{\min}(r)=r$, $l_{\max}(r)=f_{\max}$.
The sequence of operations is as follows.
  1. 1.
    Pick a random integer $u$ according to a geometric distribution $p_{u}$. As we’ll see, $u$ is the total number of operators in the ‘gaps’.
  2. 2.
    If $q<u+r_{\min}$, then the update is rejected.
  3. 3.
    Pick a random integer $r$ such that $r_{\min}\leq r\leq\min(r_{\max},q-u)$. We note that the probability $p_{r}(q)=(\min(r_{\max},q-u)-r_{\min}+1)^{-1}$ depends on $q$.
  4. 4.
    Randomly pick a sub-sequence $\widetilde{S}$ of length $r+u$ containing consecutive operators from the sequence $S_{{\bf i}_{q}}$.
  5. 5.
    Randomly choose a subsequence $S$ of length $r$ from $\widetilde{S}$. The remaining $u$ operators in $\widetilde{S}$ we will call ‘gaps’.
  6. 6.
    With probability $1/2$ we set the inv flag to true, otherwise to false.
  7. 7.
    Find all fundamental cycles of lengths $l$ such that $l_{\min}(r)\leq l\leq l_{\max}(r)$, each containing all operators of $S$ if ${\text{\it}inv}=\text{\it false}$, or all operators of $S^{-1}$ if ${\text{\it}inv}=\text{\it true}$. Here, $S^{-1}$ denotes the sequence in which each operator from $S$ is replaced by its inverse operator. Denote by $n_{c}$ the number of found cycles.
  8. 8.
    If $n_{c}=0$, the update is rejected. Otherwise, we randomly choose one of the found fundamental cycles. Let us denote by $S^{\prime}$ the sequence consisting of all the remaining $r^{\prime}$ operators from the selected cycle.
  9. 9.
    Attempt to replace the sub-sequence $\widetilde{S}$ of length $r+u$ by the sequence $\widetilde{S^{\prime}}$ of length $r^{\prime}+u$. Here, $\widetilde{S}$ contains all operators of $S^{\prime}$, as well as all the ‘gaps’ if ${\text{\it}inv}=\text{\it true}$, and otherwise it contains all operators of $S^{\prime-1}$, as well as all the ‘gaps’. We shuffle the sequence $\widetilde{S^{\prime}}$ so that its operators are contained in random order. We accept the update with the probability $P_{accept}$, which is considered below.
Compared to the spin-1/2 version of this algorithm outlined in Ref. [8], here the sequence $S$ may contain repetitions. Let the sequence $S$ contain $s_{i}$ operators $P_{i}$ such that $\sum_{i}s_{i}=r$, and the sequence $S^{\prime}$ contain $s^{\prime}_{i}$ operators $P_{i}$ such that $\sum_{i}s^{\prime}_{i}=r^{\prime}$. Let us find the acceptance probability $P_{accept}$ such that the detailed balance holds for the above protocol. Suppose that the $u$ gaps contain $u_{i}$ of operators $P_{i}$, where $i=1,2,\dots,M$, so that $\sum_{i}u_{i}=u$. Let us denote the old and new configurations as $A$ and $B$, probability to select $B$ from $A$ as $P_{select}(A\to B)$, probability to select $A$ from $B$ as $P_{select}(B\to A)$. Then, we have
$P_{select}(A\to B)=p_{u}\cdot p_{r}(q)\cdot(q-(r+u)+1)^{-1}\frac{1}{n_{c}}\times\\ \times\binom{r+u}{u}^{-1}\binom{r^{\prime}+u}{u}^{-1}\cdot\frac{s^{\prime}_{1}!\dots s^{\prime}_{M}!}{r^{\prime}!}\cdot\frac{u_{1}!\dots u_{M}!}{u!},$ (42)
$P_{select}(B\to A)=p_{u}\cdot p_{r}(q^{\prime})\cdot(q^{\prime}-(r^{\prime}+u)+1)^{-1}\frac{1}{n^{\prime}_{c}}\times\\ \times\binom{r^{\prime}+u}{u}^{-1}\binom{r+u}{u}^{-1}\cdot\frac{s_{1}!\dots s_{M}!}{r!}\cdot\frac{u_{1}!\dots u_{M}!}{u!}.$ (43)
Here, $n^{\prime}_{c}$ is the number of fundamental cycles of lengths $l$ such that $l_{\min}(r^{\prime})\leq l\leq l_{\max}(r^{\prime})$, each containing all $r^{\prime}$ operators of the sub-sequence $S^{\prime}$. Since $q^{\prime}=q+r^{\prime}-r$, we have $q-(r+u)+1=q^{\prime}-(r^{\prime}+u)+1$. Therefore,
$P_{accept}(A\to B)=\min\left(1,\frac{W_{B}}{W_{A}}\cdot\frac{P_{select}(B\to A)}{P_{select}(A\to B)}\right)$
$=\min\left(1,\frac{W_{B}}{W_{A}}\cdot\frac{p_{r}(q^{\prime})}{p_{r}(q)}\cdot\frac{n_{c}}{n^{\prime}_{c}}\cdot\frac{r^{\prime}!}{s^{\prime}_{1}\dots s^{\prime}_{M}}\cdot\frac{s_{1}!\dots s_{M}!}{r!}\right).$ (44)
Here, $W_{A}$ and $W_{B}$ are the weights of the old and the new operator sequences. Because $P(A\to B)=P_{select}(A\to B)P_{accept}(A\to B)$ and $P(B\to A)=P_{select}(B\to A)P_{accept}(B\to A)$, Eq. (44) satisfies the detailed balance condition.
Appendix B Solving systems of modular equivalences
A system of modular equivalences Eq. (35) can be rewritten as a system of Diophantine equations
$\sum_{i=1}^{M}a_{i}\cdot n_{i,j}+(2s_{j}+1)\cdot k_{j}=0,\,j=1,2,\dots,n,$ (45)
where $k_{1},k_{2},\dots,k_{N}$ are additional unknown integers. The Diophantine system (45) can be solved by employing Hermite normal form of the corresponding matrix. For details, the reader is referred to Refs. [21, 22].